40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri May 9 01:36:28 2014, rana, Summary, SUS, Optical Lever QPD Sum trends: they're almost all too weak OLtrend.png
    Reply  Mon May 12 22:42:19 2014, rana, Summary, SUS, Optical Lever QPD Sum trends: they're almost all too weak OL-FB.pngarms_140512.pdf
       Reply  Mon May 12 22:52:59 2014, Jenne, Summary, SUS, Optical Lever Filters are all different OpLevFilters.pdf
       Reply  Tue May 13 13:27:58 2014, Steve, Summary, SUS, Optical Lever Servos setting table OLsums.png
          Reply  Tue May 13 17:03:05 2014, Steve, Summary, SUS, ETMX oplev laser replaced 
             Reply  Tue May 13 17:31:32 2014, Jenne, Summary, SUS, ETMX oplev laser replaced: New oplev gains set OLerr_13May2014.pdfOLfb_13May2014.pdf
                Reply  Tue May 13 22:55:57 2014, Jenne, Summary, SUS, ETMX oplev: cleanup 
             Reply  Wed May 14 10:04:06 2014, Steve, Summary, SUS, oplev laser and temperature TempRules.png
                Reply  Tue May 20 09:12:25 2014, Steve, Summary, SUS, ETMX oplev qpd board ETMXOLQPD121.jpgETMXOLQPDface121.jpgETMXolSUM.png
                   Reply  Tue May 20 15:54:39 2014, rana, Summary, SUS, ETMX oplev qpd board 
Message ID: 9973     Entry time: Tue May 20 09:12:25 2014     In reply to: 9952     Reply to this: 9975
Author: Steve 
Type: Summary 
Category: SUS 
Subject: ETMX oplev qpd board 

Quote:

 ETMX oplev qpd gain has to be increased.

 

 Atm3, Oplev sum read 12,000 counts when the qpd was disconnected ?

           Dark qpd was zero and normal He/Ne incident on qpd was 1,730 counts.

Attachment 1: ETMXOLQPD121.jpg  585 kB  Uploaded Tue May 20 10:18:34 2014  | Hide | Hide all | Show all
ETMXOLQPD121.jpg
Attachment 2: ETMXOLQPDface121.jpg  511 kB  Uploaded Tue May 20 10:19:03 2014  | Show | Hide all | Show all
Attachment 3: ETMXolSUM.png  16 kB  Uploaded Tue May 20 13:39:47 2014  | Hide | Hide all | Show all
ETMXolSUM.png
ELOG V3.1.3-